# An Iconic Heavy Hitter Algorithm Made Private

### v1

@@ Abstract
Identifying heavy hitters in data streams is a fundamental problem with widespread applications in modern analytics systems. These streams are often derived from sensitive user activity, making update-level privacy guarantees necessary. While recent work has adapted the classical heavy hitter algorithm Misra-Gries to satisfy differential privacy in the streaming model, the privatization of other heavy hitter algorithms with better empirical utility is absent.
@@ 1 Introduction
Identifying the most frequent elements of a dataset (commonly referred to as “heavy hitters”) is a cornerstone problem in data analysis cormode2008finding. In modern systems, this task often arises on massive data streams, where items arrive continuously and must be processed online under strict memory and time constraints. Notable techniques include counter-based methods misra1982finding; mitzenmacher2012hierarchical and linear sketches charikar2002finding; cormode2005improved. However, a defining characteristic of such streams is that they are frequently derived from sensitive user activity wang2021continuous, and the classic techniques do not provide update-level privacy.
@@
Recent work seeks to retrofit these algorithms with differential privacy ($\mathsf{DP}$), aiming for minimal overhead and limited impact on utility biswas2024differentially; lebeda2023better; pagh2022improved; zhao2022differentially. At a high level, $\mathsf{DP}$ ensures that the outcome of an analysis is stable under small changes to the underlying dataset, thereby limiting what can be inferred about any single update. This line of inquiry commonly adopts the single-observation streaming model: the stream is processed in one pass and a single $\mathsf{DP}$ summary is released. This setting aligns with widely deployed analytics pipelines that aggregate large volumes of data and publish sanitized summaries under a fixed privacy budget.
@@
From an algorithmic perspective, enforcing $\mathsf{DP}$ for heavy hitters exposes two distinct challenges. First, noise must be added to protect individual updates without overwhelming the frequency signal of genuinely heavy items. Second, the algorithm must report item identities, not just noisy frequency estimates, without leaking information about rare or isolated labels. Prior work has addressed these challenges along two complementary lines. One line focuses on the counter-based method Misra-Gries ($\mathsf{MisraGries}$), which explicitly maintain candidate items and their counts biswas2024differentially; chan2012differentially; lebeda2023better. The other centers on utilizing $\mathsf{DP}$ linear sketches as frequency oracle within a (labor intensive) heavy hitter search algorithm bassily2017practical.
@@
Building on initial work by Chen et al. chan2012differentially, Lebeda and Tetek adapted the $\mathsf{MisraGries}$ algorithm for private release with asymptotically optimal noise error. This is significant as the authors note that “In practice, non-private approximate histograms are often computed using the Misra-Gries sketch”. However, empirically, the counter-based method SpaceSaving ($\mathsf{SpaceSaving}$) mitzenmacher2012hierarchical almost universally outperforms $\mathsf{MisraGries}$ (and linear sketches) in utility. This empirical dominance raises the natural question as to whether $\mathsf{SpaceSaving}$ can be adapted to satisfy $\mathsf{DP}$ in the single-observation streaming model. In response, we provide the first method to release $\mathsf{SpaceSaving}$ under $\mathsf{DP}$ and demonstrate that it retains its empirical dominance in practice.
@@
In contrast, linear sketches have been adapted to admit $\mathsf{DP}$ frequency estimates pagh2022improved; zhao2022differentially with strong utility. However, uncritically turning those estimates into released labels is not de facto differentially private. The challenge is reporting frequent items without leaking the presence or absence of isolated labels across similar streams. A naive approach is to iterate across the entire input domain, a prohibitive computational effort in most settings. To overcome this issue, Bassily et al. provide a more efficient search method for finding heavy hitters that utilizes a linear sketch as a $\mathsf{DP}$ frequency oracle bassily2017practical in post-processing. However, this approach still requires significant computational effort and memory requirements that exceed allocations that are standard in the data stream model of computation. To overcome this issue, building on observations from the $\mathsf{DP}$ release of $\mathsf{SpaceSaving}$, we provide a generic method for extracting heavy hitters from any $\mathsf{DP}$ frequency oracle at an additional memory cost of $\mathcal{O}(k)$, where $k$ is the number of heavy items in the stream.
1.1 Our contributions.
@@
In providing novel methods for privatizing existing non-private methods for identifying heavy hitters, our contributions can be summarized as follows.
  • We design the first $\mathsf{DP}$ variant of $\mathsf{SpaceSaving}$ tailored to the single-observation model. Our construction processes the output of a non-private $\mathsf{SpaceSaving}$ by injecting calibrated noise and applying a selection rule that suppresses privacy leakage events. The calibrated noise added is asymptotically optimal and the algorithm retains the performance of non-private $\mathsf{SpaceSaving}$.
  • We provide a general method to output frequent items from any $\mathsf{DP}$ frequency oracle in the single-observation model. The method wraps the oracle with a label-private selection procedure that suppresses unstable labels and releases only items whose noisy estimates exceed a $\mathsf{DP}$-calibrated threshold. The additional memory cost is proportional to the number of heavy items and can be immediately utilized with existing $\mathsf{DP}$ linear sketches.
@@
  • We perform experimental study on synthetic and real-world datasets, comparing our private heavy hitter algorithms against the state-of-the-art $\mathsf{DP}$ $\mathsf{MisraGries}$. Across a wide range of privacy parameters and space budgets, we show that $\mathsf{SpaceSaving}$ achieves superior utility while matching $\mathsf{MisraGries}$ in memory and throughput. This demonstrates the empirical advantages of $\mathsf{SpaceSaving}$ survive privatization.
1.2 Algorithm overview.
A central challenge in privatizing counter-based streaming algorithms is that the component data structures are used to track frequent items. The tracked item sets explicitly contain IDs, and these labels can leak sensitive update-level information if the output is not processed correctly. At the heart of our approach is a strategy of characterizing the differences between the tracked sets on neighboring streams, which lets us separate unstable labels from stable ones and process releases in a $\mathsf{DP}$-safe manner. Neighboring streams $X,X^{\prime}$ are defined as streams that differ in one update ($X=X^{\prime}\cup\{x\}$).
@@
Operationally, our $\mathsf{DP}$ $\mathsf{SpaceSaving}$ runs a non-private $\mathsf{SpaceSaving}$ summary with a slightly expanded capacity to facilitate isolating low-count, unstable labels. Based on our characterization of tracked set differences on neighboring streams, at release time, we add independent Laplace noise to the counters and apply a single threshold that both (i) suppresses labels that can differ across neighboring inputs and (ii) retains true heavy hitters with high probability. The expanded counter capacity allows us to set thresholds that target tracked labels at risk of breaching privacy without eliminating true heavy hitters and, thus, impeding algorithm recall.
@@ 2.1 Privacy
$\mathsf{DP}$ ensures that the inclusion or exclusion of any individual in a dataset has a minimal and bounded impact on the output of a mechanism. Two streams $\mathcal{X}=\{x_{1},\ldots,x_{T}\}$ and $\mathcal{X}^{\prime}=\{x_{1}^{\prime},\ldots,x_{T}^{\prime}\}$ are neighboring, denoted $\mathcal{X}\sim\mathcal{X}^{\prime}$, if they differ in one element. Formally, $\mathcal{X}\sim\mathcal{X}^{\prime}$ if there exists a unique $i$ such that $x_{i}\neq x_{i}^{\prime}$. The following definition of $\mathsf{DP}$ is adapted from Dwork and Roth dwork2014algorithmic.
@@
Let $\mathcal{X},\mathcal{X}^{\prime}\in\mathcal{U}^{*}$ be neighboring input streams (i.e., differing in at most one element). A randomized mechanism $\mathcal{M}$ satisfies $(\varepsilon,\delta)$-differential privacy under a given observation model if for all measurable sets of outputs $Z\subseteq\mathtt{support}(\mathcal{M})$, it holds that:
$\mathds{P}[\mathcal{M}(\mathcal{X})\in Z]\leq e^{\varepsilon}\cdot\mathds{P}[\mathcal{M}(\mathcal{X}^{\prime})\in Z]+\delta.$
In the single-observation model, the mechanism processes the stream once and releases a single $\mathsf{DP}$ output after the final stream update has been observed. In this model, neighboring streams differ by the addition or removal of a single update.
A common method for outputting $\mathsf{DP}$ functions is the following.
@@
Let $f:\mathcal{X}\to\mathbb{R}^{d}$ be a function with $\ell_{1}$-sensitivity at most $1$, that is, for any pair of neighboring inputs $X,X^{\prime}\in\mathcal{X}$,
$\|f(X)-f(X^{\prime})\|_{1}\leq 1.$
@@
where each $Z_{i}\sim\mathtt{Laplace}(1/\varepsilon)$ independently. Then $\mathcal{M}$ satisfies $(\varepsilon,0)$-$\mathsf{DP}$.
@@ 2.2 Heavy Hitter Detection
Streaming algorithms must use space sublinear in the stream lenth and universe size and process each update quickly.
@@ 3.1 Counter Methods
2 $\mathcal{T},C\leftarrow\varnothing$
@@
4    if $x_{t}\in\mathcal{T}$ then
5       $C[x_{t}]\leftarrow C[x_{t}]+1$
6    else if $|\mathcal{T}|<k$ then
7       $\mathcal{T}\leftarrow\mathcal{T}\cup\{x_{t}\}$
8       $C[x_{t}]\leftarrow 1$
@@
10       for $y\in\mathcal{T}$ do
11          $C[y]\leftarrow C[y]-1$
12          if $C[y]=0$ then
13             $\mathcal{T}\leftarrow\mathcal{T}\setminus\{y\}$
14 return $\mathcal{T},C$
@@ 13       $\mathcal{T}_{t}\leftarrow(\mathcal{T}_{t-1}\setminus\{y\})\cup\{x_{t}\}$
Counter-based methods maintain a small table of size $\mathcal{O}(k)$ of item-counter pairs. On each stream update, they adjust counters so that one obtains error guarantees on both frequency estimation and heavy-hitter recovery. These methods are attractive for their simplicity and deterministic error bounds.
@@ 3.1.1 Misra Gries
The $\mathsf{MisraGries}$ algorithm maintains at most k1k-1 $(\text{item},\text{counter})$ pairs. For each incoming item, if it is already present in the table $\mathcal{T}$, its counter is incremented. If it is not present and a counter has value zero, that counter is assigned to the new item with count set to 1. If all k1k-1 counters are occupied by distinct items, all counters are decremented by 1. A simple grouping argument shows that any item appearing more than $n/k$ times in the stream is guaranteed to be stored in the table when the algorithm terminates. The process is formalized in Algorithm 1.
To privatize a counter-based algorithm it is not sufficient to inject noise into the counts of the tracked items. This is because the IDs of the tracked items can differ on neighboring inputs, allowing an adversary to easily distinguish the streams based on the output of the algorithm. This notion is formalized in the following result by Lebeda and Tetek lebeda2023better, which describes the observed outcomes when a $\mathsf{MisraGries}$ sketch processes neighboring input streams.
Lemma 2 (lebeda2023better).
Let $X=X^{\prime}\cup\{x\}$. Let $(\mathcal{T},\mathcal{C})\leftarrow\mathtt{MG}(X,k)$ and $(\mathcal{T}^{\prime},\mathcal{C}^{\prime})\leftarrow\mathtt{MG}(X^{\prime},k)$ denote the outputs of the Misra Gries algorithm on inputs $X$ and $X^{\prime}$. Then $|\mathcal{T}\cap\mathcal{T}^{\prime}|\geq k-2$; for all $y\notin\mathcal{T}\cap\mathcal{T}^{\prime}$, $C[y]\leq 1$ and $C^{\prime}[y]\leq 1$; and exactly one of the following is true:
@@ 3.1.2 Space Saving
The $\mathsf{SpaceSaving}$ algorithm maintains $k$ counters to track frequent items. When an item $x$ arrives, if it is already present in the table, its counter is incremented. If it is not present and a free counter exists, the item is inserted with count set to 1. If all $k$ counters are occupied, the item with the minimum count $c_{\min}$ is replaced by $i$, and its counter is set to $C[x]=c_{\min}+1$. When multiple items observe the minimum count ties can be broken arbitrarily. However, to enable privatization, we enforce the rule that the item that most recently occurred is evicted. The process is formalized in Algorithm 2. This procedure guarantees that for every item, the estimated frequency satisfies the following bound.

### v2

@@ Abstract
Identifying heavy hitters in data streams is a fundamental problem with widespread applications in modern analytics systems. These streams are often derived from sensitive user activity, making adequate privacy guarantees necessary. While recent work has adapted the classical heavy hitter algorithm Misra-Gries to satisfy differential privacy in the streaming model, the privatization of other heavy hitter algorithms with better empirical utility is absent.
@@ 1 Introduction
Identifying the most frequent elements of a dataset (commonly referred to as “heavy hitters”) is a cornerstone problem in data analysis Cormode and Hadjieleftheriou (2008). In modern systems, this task commonly arises on massive data streams, where items arrive continuously and must be processed online under strict memory and time constraints. Prominent solutions include counter-based methods Misra and Gries (1982); Mitzenmacher et al. (2012) and linear sketches Charikar et al. (2002); Cormode and Muthukrishnan (2005). However, data streams are often derived from sensitive user activity Wang et al. (2021), and the classic techniques do not provide adequate privacy.
@@
Recent work seeks to retrofit these algorithms with differential privacy ($\mathsf{DP}$) guarantees, aiming for minimal overhead and limited impact on utility Biswas et al. (2024); Lebeda and Tetek (2023); Pagh and Thorup (2022); Zhao et al. (2022). At a high level, $\mathsf{DP}$ ensures that the outcome of an analysis does not change significantly under small changes to the underlying dataset, thereby limiting what can be inferred about any single update. This line of inquiry commonly adopts the single-observation streaming model: the stream is processed in one pass and a single $\mathsf{DP}$ solution is released. The setting reflects common analytics pipelines that aggregate large volumes of data and publish sanitized summaries under a fixed privacy budget.
@@
From an algorithmic perspective, enforcing $\mathsf{DP}$ for heavy hitters exposes two distinct challenges. First, sufficient noise must be added to protect individual updates without overwhelming the frequency signal of genuinely heavy items. Second, the algorithm must report item identities, even though the set of reported items may change between similar streams. Consequently, identity disclosure is a source of privacy leakage. Prior work has addressed these challenges along two complementary lines. One line focuses on the counter-based method Misra-Gries ($\mathsf{MisraGries}$), which explicitly maintains candidate frequent items and their counts Biswas et al. (2024); Chan et al. (2012); Lebeda and Tetek (2023). The other centers on utilizing $\mathsf{DP}$ linear sketches as frequency oracles within a (labor intensive) heavy hitter search algorithm Bassily et al. (2017).
@@
Building on initial work by Chen et al. Chan et al. (2012), Lebeda and Tetek adapted the $\mathsf{MisraGries}$ algorithm to enable $\mathsf{DP}$ releases with asymptotically optimal noise error Lebeda and Tetek (2023). This result is notable as the authors observe that “in practice, non-private approximate histograms are often computed using the Misra-Gries sketch”. Nevertheless, empirical studies consistently show that the counter-based SpaceSaving algorithm ($\mathsf{SpaceSaving}$) Mitzenmacher et al. (2012) substantially outperforms $\mathsf{MisraGries}$ (and linear sketches) in utility Cormode and Hadjieleftheriou (2008). This gap raises the natural question as to whether $\mathsf{SpaceSaving}$ can be adapted to satisfy $\mathsf{DP}$ in the single-observation streaming model. In response, we provide the first method to release $\mathsf{SpaceSaving}$ under $\mathsf{DP}$ and demonstrate that it retains its empirical dominance in practice.
@@
In parallel, linear sketches have been adapted to admit $\mathsf{DP}$ frequency estimates with strong utility Pagh and Thorup (2022); Zhao et al. (2022). However, as idenity disclosure is a form of privacy leakage, uncritically converting these $\mathsf{DP}$ estimates into released item identities is not, in itself, differentially private. A naive brute-force solution is to query the entire input domain, a prohibitive computational effort in most settings. To address this, Bassily et al. propose a more efficient heavy-hitter search procedure that utilizes a linear sketch as a $\mathsf{DP}$ frequency oracle in post-processing Bassily et al. (2017). Despite this improvement, the approach still incurs substantial computational and memory overheads that exceed the budgets typically assumed in the data stream model. Building on insights from our $\mathsf{DP}$ release of $\mathsf{SpaceSaving}$, we introduce a generic method for extracting heavy hitters from any $\mathsf{DP}$ frequency oracle using only an additional $\mathcal{O}(k)$ memory, where $k$ is the number of heavy items. In contrast to prior work, this $\mathcal{O}(k)$ overhead is compatible with standard streaming constraints.
1.1 Our contributions
@@
In providing novel methods for privatizing existing algorithms for identifying heavy hitters, our contributions can be summarized as follows.
  • We present the first $\mathsf{DP}$ variant of the $\mathsf{SpaceSaving}$ algorithm in the single-observation streaming model, achieving asymptotically optimal noise while preserving the empirical performance advantages of non-private $\mathsf{SpaceSaving}$.
  • We introduce a general framework for releasing heavy hitters from any $\mathsf{DP}$ frequency oracle in the single-observation model, with an additional memory overhead of $\mathcal{O}(k)$, where $k$ is the maximum number of heavy items.
@@
  • We conduct an experimental evaluation on synthetic and real-world datasets, comparing our methods against state-of-the-art $\mathsf{DP}$ $\mathsf{MisraGries}$. Across a wide range of privacy parameters and space budgets, our $\mathsf{DP}$ $\mathsf{SpaceSaving}$ consistently achieves higher utility while matching $\mathsf{DP}$ $\mathsf{MisraGries}$ in memory usage and throughput.
1.2 Algorithm overview
Counter-based streaming algorithms explicitly maintain sets of candidate frequent item identities. As these sets may differ between similar streams, the disclosure of item identities itself becomes a source of privacy leakage. Thus, a central challenge in privatizing a counter-based algorithm is designing a post-processing mechanism that releases frequent item sets without exposing the privacy leakage.
Neighboring streams $X,X^{\prime}$ are defined as streams that differ in one update ($X=X^{\prime}\cup\{x\}$). At the heart of our approach is a strategy of characterizing how tracked sets change across neighboring streams. This allows us to distinguish stable labels from unstable ones and to release only the former in a $\mathsf{DP}$-safe manner. Our characterization states that unstable labels always observe a frequency estimate close to the minimum estimate across tracked items. Thus, unstable labels can be suppressed with a sufficiently high threshold.
@@ 1.2 Algorithm overview
Operationally, our $\mathsf{DP}$ $\mathsf{SpaceSaving}$ runs a non-private $\mathsf{SpaceSaving}$ algorithm with a slightly expanded capacity. At release time, we add independent Laplace noise to the counters and apply a single threshold that both (i) suppresses labels that can differ across neighboring inputs and (ii) retains true heavy hitters with high probability. The expanded counter capacity allows us to reduce the threshold and, therefore, target unstable labels without eliminating true stable heavy hitters.
@@ 2.1 Privacy
$\mathsf{DP}$ ensures that the inclusion or exclusion of any individual in a dataset has a minimal and bounded impact on the output of a mechanism. Two streams ${X}=\{x_{1},\ldots,x_{T}\}$ and ${X}^{\prime}=\{x_{1}^{\prime},\ldots,x_{T}^{\prime}\}$ are neighboring, denoted ${X}\sim{X}^{\prime}$, if they differ by the addition or removal of a single update. Formally, ${X}\sim{X}^{\prime}$ if $X=X^{\prime}\cup\{x\}$. The following definition of $\mathsf{DP}$ is adapted from Dwork and Roth Dwork et al. (2014).
@@
Let ${X},{X}^{\prime}\in\mathcal{U}^{*}$ be neighboring input streams (i.e., differing in at most one element). A randomized mechanism $\mathcal{M}$ satisfies $(\varepsilon,\delta)$-differential privacy under a given observation model if for all measurable sets of outputs $Z\subseteq\mathtt{support}(\mathcal{M})$, it holds that:
$\mathds{P}[\mathcal{M}({X})\in Z]\leq e^{\varepsilon}\cdot\mathds{P}[\mathcal{M}({X}^{\prime})\in Z]+\delta.$
In the single-observation model, the mechanism processes the stream once and releases a single $\mathsf{DP}$ output after the final stream update has been observed.
A common method for outputting $\mathsf{DP}$ functions is the following
@@
Let $f:\mathcal{X}\to\mathbb{R}^{d}$ be a function with $\ell_{1}$-sensitivity at most $s$, that is, for any pair of neighboring inputs $X,X^{\prime}\in\mathcal{X}$,
$\|f(X)-f(X^{\prime})\|_{1}\leq s.$
@@
where each $Z_{i}\sim\mathtt{Laplace}(s/\varepsilon)$ independently. Then $\mathcal{M}$ satisfies $(\varepsilon,0)$-$\mathsf{DP}$.
@@ 3 Prior Methods
We now introduce background on counter-based methods and linear sketches. A broader discussion of prior work is provided in Appendix A.
@@ 3.1 Counter Methods
2 $\mathcal{T}_{0},C_{0}\leftarrow\varnothing$
@@
4    if $x_{t}\in\mathcal{T}_{t-1}$ then
5       $C_{t}[x_{t}]\leftarrow C_{t-1}[x_{t}]+1$
6    else if $|\mathcal{T}_{t-1}|<k$ then
7       $\mathcal{T}_{t}\leftarrow\mathcal{T}_{t-1}\cup\{x_{t}\}$
8       $C_{t}[x_{t}]\leftarrow 1$
9
@@ 8       $C_{t}[x_{t}]\leftarrow 1$
11       for $y\in\mathcal{T}_{t-1}$ do
12          $C_{t}[y]\leftarrow C_{t-1}[y]-1$
13
14       $S\leftarrow\{y\mid C_{t}[y]=0\}$
15       $\mathcal{T}_{t}\leftarrow\mathcal{T}_{t-1}\setminus S$
16
17 return $\mathcal{T}_{T},C_{T}$
@@ 13       $\mathcal{T}_{t}\leftarrow(\mathcal{T}_{t-1}\setminus\{y\})\cup\{x_{t}\}$
Counter-based methods maintain a small table of $\mathcal{O}(k)$ (item,counter) pairs. On each stream update, these methods adjust tracked items and corresponding counters so that one obtains error guarantees on both frequency estimation and heavy-hitter recovery. These methods are attractive for their simplicity and deterministic error bounds.
@@ 3.1.1 Misra Gries
The $\mathsf{MisraGries}$ algorithm maintains a set $\mathcal{T}$ and corresponding counters $C$, both of size at most k1k-1. For each incoming item, if it is already present in the table $\mathcal{T}$, its counter is incremented. If it is not present and a counter has value zero, that counter is assigned to the new item with count set to 1. If all k1k-1 counters are occupied by distinct items, all counters are decremented by 1. For a stream of length $T$, a simple grouping argument shows that any item appearing more than $T/k$ times in the stream is guaranteed to be stored in the table when the algorithm terminates. The process is formalized in Algorithm 1.
To privatize a counter-based algorithm it is not sufficient to inject noise into the counts of the tracked items. This is because the identities of the tracked items can differ on neighboring inputs, allowing an adversary to easily distinguish the streams based on the output of the algorithm. This notion is formalized in the following result by Lebeda and Tetek Lebeda and Tetek (2023), which describes the observed outcomes when a $\mathsf{MisraGries}$ algorithm processes neighboring input streams.
Lemma 2 (Lebeda and Tetek (2023)).
Let $X=X^{\prime}\cup\{x\}$. Let $(\mathcal{T},{C})\leftarrow\mathtt{MG}(X,k)$ and $(\mathcal{T}^{\prime},{C}^{\prime})\leftarrow\mathtt{MG}(X^{\prime},k)$ denote the outputs of the Misra Gries algorithm on inputs $X$ and $X^{\prime}$. Then $|\mathcal{T}\cap\mathcal{T}^{\prime}|\geq k-2$; for all $y\notin\mathcal{T}\cap\mathcal{T}^{\prime}$, $C[y]\leq 1$ and $C^{\prime}[y]\leq 1$; and exactly one of the following is true:
@@ 3.1.2 Space Saving
The $\mathsf{SpaceSaving}$ algorithm has the same component data structures ($\mathcal{T},C$) as $\mathsf{MisraGries}$. When an item $x$ arrives, if it is already present in the table, its counter is incremented. If it is not present and a free counter exists, the item is inserted with count set to 1. In contrast to $\mathsf{MisraGries}$, if all $k$ counters are occupied, the item with the minimum count $c_{\min}$ is replaced by $i$, and its counter is set to $C[x]=c_{\min}+1$. When multiple items observe the minimum count ties can be broken arbitrarily. However, to enable privatization, we enforce the rule that the item that most recently occurred is evicted. The process is formalized in Algorithm 2. This procedure guarantees that for every item, the estimated frequency satisfies the following bound.
@@
While $\mathsf{SpaceSaving}$ has the same absolute error bound as $\mathsf{MisraGries}$, empirical studies show that it has superior utility in practice Cormode and Hadjieleftheriou (2008). Currently, no $\mathsf{DP}$ version of $\mathsf{SpaceSaving}$ exists. This is the main gap addressed by our paper.